Status Report

NOAA GOES SATOPS Morning Report 20 November 2000

By SpaceRef Editor
November 20, 2000
Filed under ,



The NOAA-14 XSU was swapped on November 19.  Details below. 
No change in the status of the other POES spacecraft.

Engineering was called in on Friday night at 0020z to investigate HRPT
and stored playback problems on NOAA-14.  The XSU1 PLL2 signal, normally
steady at 4.92 volts, was history trended and it was shown that it had
sustained periods of instability over the past 24 hours, with some subtle
changes noted as early as ~day 275.  Data retrieval efforts on Saturday
continued to degrade and the decision was made to swap XSU power supply/frequency
synthesizer sides after the blind at 0955z (rev 30351) on Sunday morning. 
The swap occurred without incident and the new unit, which has not been
activated since before launch (December 1994), appears to be performing

Note:  NOAA-14 data recovery figures will not be published in this
report, as before 0955z on Sunday, they were significantly degraded across
the board before the XSU swap.  Data since the swap has been almost
perfect, except for a noisy G2B (94.6% recovered, with no re-ingest due
recorder scheduling) on rev 30357 / W at 1932z on November 19.

On Friday, November 17 (rev 13071 / F at 1720z), the NOAA-15 HIRS filter
wheel was cycled off/on  and the Filter Wheel Housing Heater was turned
back on for the weekend as part of the continuing NOAA-15 HIRS Filter Wheel
Test (Procedure # 001115.1525).  Additional actions this week include
MACRO toggling at the FW low power mode.

A NOAA-16 APT VTX2 anomaly telecon has been scheduled for today at 1300

It is anticipated that the NOAA-15 AVHRR anomaly testing will be coordinated
and approved this week for execution next week after the holiday.

POES operations were otherwise nominal over the past 72 hours.


No change in the status of the DMSP spacecraft.

DMSP operations were nominal over the past 72 hours.


No change in the status of the GOES spacecraft.

At 1731z on Sunday, November 19, a GOES-8 frame break was observed and
attributed to high CRC errors.

GOES-9 and GOES-11 remain in the Z-Axis Precession (ZAP) storage mode
at ~105 deg W.

GOES operations were nominal over the past 72 hours.






Questions or comments to

SpaceRef staff editor.